OpenStudy (anonymous):

A cruise ship has circular windows in its state rooms. What is the circumference of one of these windows If it's diameter is 15 inches??

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

C = pi*d = 15pi

OpenStudy (anonymous):

so.. 3.14(15) ?

OpenStudy (skullpatrol):

@t0mmiere could you please provide some explanation?

OpenStudy (skullpatrol):

The number $$\pi$$is defined as $$\frac{C}{d}$$where C is the circumference and d is the diameter.

OpenStudy (skullpatrol):

Any questions so far @xoxox123 ?

OpenStudy (anonymous):

ok i get that

OpenStudy (skullpatrol):

$$\pi=\frac{C}{d}$$ $$C=\pi*d$$

OpenStudy (anonymous):

c= 3.14(15)???

OpenStudy (skullpatrol):